In a circle O, chords AB and CD intersect at E (not center).
AE = X, EB = X+1, CE = X-1 and ED = 2X.
Find AE.

Answer by venugopalramana(3286)

In a circle O, chords AB and CD intersect at E (not center).
AE = X, EB = X+1, CE = X-1 and ED = 2X.
Find AE.
IN A CYCLIC QUADRILATERAL,WE HAVE
AE*EB=CE*ED
X(X+1)=(X-1)2X
X+1=2X-2
2X-X=1+2=3
X=3
SO..
AE=X=3